OpenStudy (anonymous):

Can someone help me out on multiplying and simplifying x-6/x^2 + 4x - 32 * x-4/x+2 ?

OpenStudy (anonymous):

\[\large \frac{x-6}{x^2+4x-32}\times \frac{x-4}{x+2}\]?

OpenStudy (anonymous):

yes!

OpenStudy (anonymous):

not much to do maybe we can factor the denominator and cancel something

OpenStudy (anonymous):

yeah i think that's what it calls for, but i'm so confused on how to do any of that

OpenStudy (anonymous):

\[x^2+4x-32=(x+8) (x-4)\]

OpenStudy (anonymous):

?

OpenStudy (anonymous):

so you do get to cancel one factor but that is it \[\large \frac{x-6}{x^2+4x-32}\times \frac{x-4}{x+2}\\ =\frac{(x-6)(x-4)}{(x+8)(x-4)(x+2)}\\ =\frac{(x-6)}{(x+8)(x+2)}\]

OpenStudy (anonymous):

cancelled the common factor top and bottom of \(x-4\)

OpenStudy (anonymous):

OHHHH, this i understand! :D the notes i have on it are so confusing. thank you so much, can you help me with two more? or explain the rule for it?

OpenStudy (anonymous):

sure you want to post it here?

OpenStudy (anonymous):

x^2 - 4/x^2 - 1 * x+1/x+2

OpenStudy (anonymous):

\[\frac{x^2-4}{x^2-1}\times \frac{x+1}{x+2}\] again factor and cancel \[\large \frac{(x+2)(x-2)(x+1)}{(x+1)(x-1)(x+2)}\]

OpenStudy (anonymous):

this time you can cancel two factors clear?

OpenStudy (anonymous):

yeah i think so

OpenStudy (anonymous):

both factoring was from the difference of two square \[a^2-b^2=(a+b)(a-b)\] so \[x^2-4=(x+2)(x-2)\] in the top and \[x^2-1=(x+1)(x-1)\] in the bottom

OpenStudy (anonymous):

you are left with \[\large \frac{\cancel{(x+2)}(x-2)\cancel{(x+1)}}{\cancel{(x+1)}(x-1)\cancel{(x+2)}}\]\[=\frac{x-2}{x-1}\]

OpenStudy (anonymous):

I think I'm getting it, wow I'm surprised! Can you help me with just this last one then I'll stop bugging you?

OpenStudy (anonymous):

k no problem

OpenStudy (anonymous):

x+3/x * x+4/x^2 + 7x + 12

OpenStudy (anonymous):

\[\frac{x+3}{x}\times \frac{x+4}{x^2+7x+12}\]?

OpenStudy (anonymous):

yeah, this one is a bit confusing :/

OpenStudy (anonymous):

this one you can cancel your brains out because \[x^2+7x+12=(x+3)(x+4)\]

OpenStudy (anonymous):

\[\frac{x+3}{x}\times \frac{x+4}{x^2+7x+12}\]\[=\frac{x+3}{x}\times \frac{x+4}{(x+3)(x+4)}\]

OpenStudy (anonymous):

now everything goes leaves only \(1\) in the top and \(x\) in the bottom

OpenStudy (anonymous):

\[=\frac{(x+3)(x+4)}{x(x+3)(x+4)}\] \[=\frac{1}{x}\]

OpenStudy (anonymous):

I think I am proud to say I fully understand it now! It took me forever by just looking at my textbook. Thanks for the help :)

OpenStudy (anonymous):

yw glad to help good luck!